OpenStudy (anonymous):

Evaluate the expression 4 × z + 12 when z = 15.

OpenStudy (anonymous):

72

OpenStudy (anonymous):

\[4 \times z + 12\] Since z = 15, we can substitute z for 15. \[4\times 15 + 12\] \[60 + 15\] \[72\]